The capital cost of Waste-Water Treatment Plant for Gambang Town of 150,000 people
was estimated to be RM75.00 per person in 1995. If the modest estimate of the rate of
inflation is 4.0% for the period to 2015, what is the per capita capital cost of the
Treatment Plant now?
Answer #1

for 1995 capital cost is 75 per person.

so total capital cost for Waste-Water Treatment Plant= 75*150000

=11250000

here rate of infletion is given. means every year 4% will cumpounded in capital cost.

A=P(1+r)n

A=11250000(1+0.04)20

(r= interest rate, n= cumpound year time duration, P= amount)

A=24650135

so total capital cost for Waste-Water Treatment Plant is 24650135

so per capita capital cost of theTreatment Plant = 24650135/ total population

=24650135/150000

=164.33